A new drug cures 9 of 200 patients suffering from a type of cancer, for which the historical cure rate is 2%. Perform a test to check on the significance of this result, at both 5% and 1% levels of significance. Based on your conclusion, comments on the efficacy of the treatment using this new drug.

(15 marks


1
Expert's answer
2021-03-17T12:29:54-0400

"H_0: p=0.02\\
H_1: p\neq 0.02\\
N=200\\
\overline{p}=\frac{9}{200}\\
\alpha_1=0.05\\
\text{We will use the following test statistic:}\\
Z=\frac{\overline{p}-p}{\sqrt{\frac{p(1-p)}{n}}}\\
\text{Our text statistic }z=\frac{9/200-0.02}{\sqrt{\frac{(0.02)(0.98)}{200}}}\approx 2.53.\\
\text{Critical value:}\\
\Phi(z_{cr})=\frac{1-\alpha_1}{2}\\
\Phi(z_{cr})=0.475\\
z_{cr}=1.96\\
(-\infty,-1.96)\cup (1.96,\infty)\text{ is the rejection region}.\\
\text{Our test statistic z falls into the rejection region. So we reject }H_0: p=0.02\\
\text{and accept }H_1: p\neq 0.02.\\
\text{The efficacy of the treatment using this new drug is not equal to 2\%}.\\
\alpha_2=0.01\\
\Phi(z_{cr})=\frac{1-\alpha_2}{2}\\
\Phi(z_{cr})=0.495\\
z_{cr}=2.58\\
(-\infty,-2.58)\cup (2.58,\infty)\text{ is the rejection region}.\\
\text{Our test statistic z does not fall into the rejection region. So we accept }\\
H_0: p=0.02.\\
\text{The efficacy of the treatment using this new drug is equal to 2\%}."